# Hierarchical Relation-augmented Representation Generalization for Few-shot Action Recognition

### v1

@@ 3.5. Training and Testing with HR2G-shot
Table 2. The impact and parameters of core components on SSv2-small (Goyal et al., 2017) and HMDB51 (Kuehne et al., 2011) under the 5-way 1-shot setting (see §4.4).
@@
Baseline $71.0$M $55.0$ $81.2$
[…]
IKT only $74.5$M (+3.5) $57.8$ $83.2$
[…]
4. Experiments
4.1. Experimental Setup
Dataset. We evaluate our HR2G-shot on five commonly used datasets, i.e., Kinetics (Carreira and Zisserman, 2017), SSv2-small (Goyal et al., 2017), SSv2-full (Goyal et al., 2017), HMDB51 (Kuehne et al., 2011), and UCF101 (Soomro et al., 2012). For HMDB51 and UCF101, we adopt the few-shot split from (Zhang et al., 2020; Bishay et al., 2019), with $31$/$10$/$10$ classes and $70$/$10$/$21$ classes for train/val/test, respectively. Following previous works (Wang et al., 2023a; Xia et al., 2023; Cao et al., 2022), we divide Kinetics, SSv2-small and SSv2-full into three splits: $64$/$12$/$24$ classes used for train, val, and test, respectively. The difference between SSv2-small and SSv2-full is that SSv2-full contains more videos per class in train set.
Evaluation. Following standard evaluation protocols (Cao et al., 2020; Xing et al., 2023c; Wang et al., 2023a; Liu et al., 2023; Yu et al., 2023), $5$-way $1$-shot and $5$-shot accuracy over $10,000$ tasks are used for evaluation.
4.2. Implementation Details
Network Architecture. HR2G-shot adopts pre-trained CLIP ViT-B (Radford et al., 2021) as our backbone for parameter-efficient fine-tuning, for a fair comparison with previous methods (Qu et al., 2024; Wang et al., 2024b; Xing et al., 2024). In IKT, we empirically set the size of temporal knowledge bank $G\!=\!40$ (see Fig. 4). By default, the number of temporal prototypes $M$ is set to $3$ (see Table 4(c)). For other hyper-parameters, we empirically set retrieval ratio $\kappa$ and momentum $\mu$ are to $0.8$ (see Table 4(b)) and $0.99$ (see Table 4(d)), respectively.
Network Training. In line with previous works (Cao et al., 2020; Huang et al., 2022; Wang et al., 2024b; Xing et al., 2023c), we uniformly and sparsely sample $T\!=\!8$ frames from each video to encode frame-level representations. During the training phase, random horizontal flipping and color jitter are adopted for data augmentation. We freeze both the CLIP visual encoder and text encoder, and only finetune lightweight adapters in the visual encoder as (Xing et al., 2024). We conducted training using the Adam (Kingma and Ba, 2014) optimizer with the multi-step scheduler. Due to limited space, more implementation details are left in the appendix.
4.3. Comparison with State-of-the-Arts
Table 1 illustrates our compelling results over the top-leading FSAR solutions on five datasets (i.e., SSv2-small (Goyal et al., 2017), SSv2-full (Goyal et al., 2017), HMDB51 (Kuehne et al., 2011), UCF101 (Soomro et al., 2012), and Kinetics (Carreira and Zisserman, 2017)). Our method reports the results on CLIP-ViT-B visual encoder. For spatial-related datasets, our method yields remarkable performance on most task settings across HMDB51, UCF101, and Kinetics. Especially for 1-shot tasks, it surpasses the previous SOTA (i.e., MA-FSAR (Xing et al., 2024)) by 2.2% on HMDB51 and 0.8% on UCF101, respectively. Temporal-related datasets (SSv2-small and SSv2-full) require models to comprehend complex temporal information, making it much more challenging. Our approach still achieves dominant results on SSv2-small and SSv2-full, surpassing other competitors across all metric task settings. This reinforces our belief that our method makes use of the temporal knowledge bank, which collects temporal cues from historical tasks, to yield discriminative temporal features. All of the above improvements across all datasets and task settings show our HR2G-shot has strong generalization for different scenes. We attribute this to the fact that we learn task-specific temporal features by transferring knowledge at both video-level and task-level.
Table 3. Ablations of different inter-video interaction manner in ISC on SSv2-small (Goyal et al., 2017) and HMDB51 (Kuehne et al., 2011) under the 5-way 1-shot setting (see §4.4).
@@ 4.3. Comparison with State-of-the-Arts
Table 4. A set of ablation studies on SSv2-small (Goyal et al., 2017) and HMDB51 (Kuehne et al., 2011) under the $5$-way $1$-shot setting (§4.4). The adopted network designs are marked in red.
  Method SSv2-small HMDB51
@@
(a) Masked Interaction Strategy
@@
(b) Retrieval Ratio
@@
(c) Prototype Number
@@
(d) Momentum Coefficient
4.4. Diagnostic Experiment
Key Component Analysis. We first investigate the effectiveness of each component in HR2G-shot, i.e., Inter-video Semantic Correlation (ISC) and Inter-task Knowledge Transfer (IKT), which is summarized in Table 2. First, our proposed ISC leads to $3.7\%$ and $2.8\%$ performance gains against the baseline on SSv2-small (Goyal et al., 2017) and HMDB51 (Kuehne et al., 2011), respectively, demonstrating the value of inter-video relation modeling. Second, after incorporating our proposed IKT into the baseline, our method improves on the two datasets by $2.8\%$ and $2.0\%$, verifying that capturing task-shared temporal knowledge by exploring inter-task relationships can yield task-specific video representations. Third, our full model HR2G-shot achieves the best performance, confirming the joint effectiveness of our overall algorithm design. Different from reported learnable model parameters in Table 5, we provide full model parameters in Table 2. As seen, our algorithm brings a modest amount of extra parameters, while leveraging such a performance leap.
Imapct of Inter-video Interaction Manner in ISC. We study the impact of our fine-grained inter-video interaction (Eq. 1) by contrasting it with video-level global attention (Wang et al., 2022) and frame-level dense attention. “video-level global attention” refers to conducting inter-video interaction via global video features, and “frame-level dense attention” means one frame token interacts with all other frame tokens in current task. As outlined in Table 3, our fine-grained inter-video interaction proves to be effective–it outperforms video-level global attention and frame-level dense attention by $3.7\%$ and $1.3\%$ on SSv2-small respectively, and efficient–its FLOPs are significantly less than frame-level dense attention. The reason why our fine-grained interaction outperforms frame-level dense attention may be that frame-level dense attention often introduces redundant or noisy interactions. Thus, we conclude that our fine-grained inter-video interaction achieves a trade-off between effectiveness and efficiency for capturing task-specific temporal cues.
@@ 4.4. Diagnostic Experiment
Figure 4. The impact of temporal bank size $G$ in IKT on SSv2-small (Goyal et al., 2017) under the $5$-way $1$-shot setting (§4.4).
Impact of Different Masked Interaction Strategies in ISC. We further analyze the influence of our masked interaction strategy on SSv2-small (Goyal et al., 2017) and HMDB51 (Kuehne et al., 2011). Fig. 3(b)(c)(d) shows three other alternative masked strategies, i.e., support-support interaction, query-support interaction and full interaction. As reported in Table 4(a), our masked interaction strategy (i.e., adaptive Intraction Mask) yields significant performance advancements compared with other alternatives. This suggests that designing proper masked interactions by respecting the difference between support and query videos contributes to better fine-grained cross-video interactions.
Impact of Different Retrieval Ratio $\kappa$ in IKT. We next investigate the impact of retrieval ratio $\kappa$ (Eq. 3), which controls the number of retrieved memory representations from the knowledge bank. Table 4(b) provides comparison results with regrad to different retrieval ratios under the 1-shot setting on SSv2-small (Goyal et al., 2017) and HMDB51 (Kuehne et al., 2011). We can clearly observe that, our algorithm performs best with a relatively large retrieval ratio (i.e., $\kappa=0.7$) on SSv2-small. This verifies that we discard the semantic-unrelated temporal knowledge in the bank to alleviate the negative effects of meaningless semantic information. When $\kappa$ is too small, the performance degrades. We speculate this is because our algorithm discards too many useful memory representations related to the current task.
@@
Figure 5. Similarity visualization between query samples ($q_{n}$) and support prototypes ($s_{n}$) with different methods in a meta-test episode from HMDB51 (Kuehne et al., 2011) (see §4.5). A higher score indicates a greater degree of similarity. The green box indicates correct prediction and the red box indicates incorrect prediction.
Impact of Temporal Prototype Number $M$ in IKT. Table 4(c) reports the performance of our HR2G-shot with regard to the number of temporal prototypes. $M\!=\!1$ means that directly treating one video as a global representation. For $M=8$, we maintain original frame-level features as temporal prototypes. After constructing learnable prototypes to summarize discriminative temporal representations, we observe consistent improvements. For example, our HR2G-shot gains the best performance (i.e., 60.2% accuracy on SSv2-small dataset) when $M\!=\!3$.
Impact of Momentum Coefficient $\mu$ in IKT. We probes the impact of momentum coefficient $\mu$ (Eq.6) on SSv2-small (Goyal et al., 2017) and HMDB51 (Kuehne et al., 2011) in Table 4(d). We can clearly observe that, our algorithm performs better with a relatively large coefficient (i.e., $\mu=0.99$), but not too large (i.e., $\mu=0.999$). When $\mu$ is too small, the performance decreases. In particular, at the extreme of no momentum (i.e., $\mu=0$), the performance significantly degrades.
Figure 6. T-SNE feature visualization of six classes learned by baseline and HR2G-shot on Kinetics (Carreira and Zisserman, 2017) and HMDB51 (Kuehne et al., 2011) ( §4.5).
[…]
Efficiency Analysis. To analyze the effectiveness of our HR2G-shot, we list comparison results with CLIP-FSAR (Wang et al., 2024b) in terms of trainable parameters, GPU memory, and inference speed in Table 5. We choose ViT-B as our visual encoder. Note that different from full model parameters in Table 2, Table 5 reports trainable parameter comparison. As shown in Table 5, compared to CLIP-FSAR, our method introduces fewer trainable parameters. Though efficient in terms of parameters, our algorithm brings slight inference delay and memory costs. However, the hierarchical relation modeling allows our method to outperform CLIP-FSAR in terms of classification accuracy, i.e., 8.5% and 5.6% accuracy improvements on HMDB51 (Kuehne et al., 2011) and SSv2-small (Goyal et al., 2017) over CLIP-FSAR, respectively.
@@
Table 5. Complexity analysis for $5$-way $1$-shot HMDB51 (Kuehne et al., 2011) and SSv2-small (Goyal et al., 2017) evaluation. Note that there are $1$ query samples per class. Here, we report trainable parameters, GPU memory, and Speed for each model. “Acc1” and “Acc2” are the accuarcy on HMDB51 and SSv2-small, respectively (§4.4).
@@
CLIP-FSAR $89.5$M $14.2$G $36.7$ms $77.1$ $54.6$
HR2G-shot(ours) $19.9$M $14.3$G $39.9$ms $85.6$ $60.2$

### v3

@@ 4.2 Implementation Details
Table 4: The impact and parameters of core components on SSv2-small [71] and HMDB51 [64] under the 5-way 1-shot setting (see §4.4.1).
@@
Baseline 100.8100.8M $55.0$ $81.2$
grayblack ISC only 103.0103.0M  (+2.2) $58.7$ $84.0$
IKT only 104.3104.3M (+3.5) $57.8$ $83.2$
grayblack HR2G-shot (Ours) 106.5106.5M  (+5.7) $\mathbf{60.2}$ $\mathbf{85.6}$
4.3 Comparison with State-of-the-Arts
Table 1, 2, and 3 illustrate our compelling results over the top-leading FSAR solutions on five datasets (i.e., SSv2-small [71], SSv2-full [71], HMDB51 [64], UCF101 [65], and Kinetics [70]). Our method reports the results on CLIP-ViT-B visual encoder. For spatial-related datasets, our method yields remarkable performance on most task settings across HMDB51, UCF101, and Kinetics. Especially for 1-shot tasks, it surpasses the previous SOTA (i.e., MA-FSAR [58]) by 2.2% on HMDB51 and 0.8% on UCF101, respectively. Temporal-related datasets (SSv2-small and SSv2-full) require models to comprehend complex temporal information, making it much more challenging. Our approach still achieves dominant results, surpassing other competitors across all metric task settings. This reinforces our belief that our method makes use of the temporal knowledge bank, which collects temporal cues from historical tasks, to yield discriminative temporal features. All of the above improvements across all datasets and task settings show our HR2G-shot has strong generalization for different scenes. We attribute this to the fact that we learn task-specific temporal features by transferring knowledge at both the video-level and task-level.
4.4 Diagnostic Experiment
Table 5: Ablation study on different inter-video interaction manners in ISC on SSv2-small [71] and HMDB51 [64] under the 5-way 1-shot setting (see §4.4.2).
@@ 4.4 Diagnostic Experiment
Table 6: Ablation study on different masked interaction strategies under the 55-way 11-shot setting on SSv2-small [71] and HMDB51 [64] (see §4.4.3).
    Masked Interaction Strategy   SSv2-small   HMDB51
@@
4.4.1 Key Component Analysis
We first investigate the effectiveness of each component in HR2G-shot, i.e., Inter-video Semantic Correlation (ISC) and Inter-task Knowledge Transfer (IKT), which is summarized in Table 4.2. First, our proposed ISC leads to $3.7\%$ and $2.8\%$ performance gains against the baseline on SSv2-small [71] and HMDB51 [64], respectively, demonstrating the value of inter-video relation modeling. Second, after incorporating our proposed IKT into the baseline, our method improves on the two datasets by $2.8\%$ and $2.0\%$, verifying that capturing task-shared temporal knowledge by exploring inter-task relationships can yield task-specific video representations. Third, our full model HR2G-shot achieves the best performance, confirming the joint effectiveness of our overall algorithm design. Different from the reported learnable model parameters in Table 11, we provide full model parameters in Table 4.2. As seen, our algorithm brings a modest amount of extra parameters, while leveraging such a performance leap.
4.4.2 Impact of Inter-video Interaction Manner in ISC
We study the impact of our fine-grained inter-video interaction (Eq. 1) by contrasting it with video-level global attention [13] and frame-level dense attention. “video-level global attention” refers to conducting inter-video interaction via global video features, and “frame-level dense attention” means one frame token interacts with all other frame tokens in the current task. As outlined in Table 5, our fine-grained inter-video interaction proves to be effective–it outperforms video-level global attention and frame-level dense attention by $3.7\%$ and $1.3\%$ on SSv2-small respectively, and efficient–its FLOPs are significantly less than frame-level dense attention. The reason why our fine-grained interaction outperforms frame-level dense attention may be that frame-level dense attention often introduces redundant or noisy interactions. Thus, we conclude that our fine-grained inter-video interaction achieves a trade-off between effectiveness and efficiency for capturing task-specific temporal cues.
Table 7: Impact of different retrieval ratio $\kappa$ in IKT under the 55-way 11-shot setting on SSv2-small [71] and HMDB51 [64] (see §4.4.4). The adopted design is marked in red.
@@ 4.4.2 Impact of Inter-video Interaction Manner in ISC
Table 8: Impact of temporal prototype number $M$ in IKT under the 55-way 11-shot setting on SSv2-small [71] and HMDB51 [64] (see §4.4.5). The adopted design is marked in red.
@@
4.4.3 Impact of Different Masked Interaction Strategies in ISC
We further analyze the influence of our masked interaction strategy on SSv2-small [71] and HMDB51 [64]. Fig. 3(b)(c)(d) shows three other alternative masked strategies, i.e., support-support interaction, query-support interaction and full interaction. As reported in Table 6, our masked interaction strategy (i.e., adaptive Intraction Mask) yields significant performance advancements compared with other alternatives. Notably, on the challenging SSv2-small benchmark, our Adaptive Interaction Mask surpasses the second-best strategy (query-support) by 1.2% and the full interaction baseline by 2.4%. It is worth noting that the full interaction strategy yields suboptimal results, indicating that simply aggregating all token interactions leads to information redundancy. Our approach addresses this by dynamically tailoring the interaction mask, thereby facilitating more precise cross-video reasoning.
4.4.4 Impact of Different Retrieval Ratio $\kappa$ in IKT
We next investigate the impact of retrieval ratio $\kappa$ (Eq. 3), which controls the number of retrieved memory representations from the knowledge bank. Table 7 provides comparison results with regard to different retrieval ratios under the 5-way 1-shot setting on SSv2-small [71] and HMDB51 [64]. We can clearly observe that, our algorithm performs best with a relatively large retrieval ratio (i.e., $\kappa=0.7$). This verifies that we discard the semantically unrelated temporal knowledge in the bank to alleviate the negative effects of meaningless semantic information. Specifically, the accuracy improves as $\kappa$ increases from $0.3$ to $0.7$, achieving the optimal performance at $\kappa=0.7$ on both benchmarks. This improvement suggests that a broader retrieval scope allows the model to access sufficient diverse temporal patterns for effective generalization. However, a further increase in $\kappa$ (to $0.9$ or $1.0$) results in a performance decline. This phenomenon verifies our hypothesis that the knowledge bank contains inevitable semantic noise or task-irrelevant temporal dynamics. Therefore, $\kappa=0.7$ strikes an optimal balance between informativeness and robustness.
Table 9: Impact of momentum coefficient $\mu$ in IKT under the 55-way 11-shot setting on SSv2-small [71] and HMDB51 [64] (see §4.4.6). The adopted design is marked in red.
@@ 4.4.4 Impact of Different Retrieval Ratio $\kappa$ in IKT
4.4.5 Impact of Temporal Prototype Number $M$ in IKT
Table 8 reports the performance of our HR2G-shot with regard to the number of temporal prototypes, which determines the granularity of the summarized temporal representations. The case of $M=1$ means that directly treating one video as a global representation. The suboptimal performance ($59.0\%$ on SSv2-small [71]) suggests that global pooling overlooks critical fine-grained temporal cues for action recognition. Our algorithm performs best with $M=3$ temporal prototypes. We hypothesize that $M=3$ aligns well with the intrinsic structure of actions (typically consisting of onset, apex, and offset phases), thus effectively summarizing discriminative patterns. However, further increasing $M$ to $8$ causes a performance drop. This indicates that maintaining excessive temporal granularity (similar to raw frame-level features) introduces redundancy and noise, hampering the generalization in the few-shot regime.
4.4.6 Impact of Momentum Coefficient $\mu$ in IKT
We probe the impact of momentum coefficient $\mu$ (Eq. 6) on SSv2-small [71] and HMDB51 [64], which regulates the update rate of the memory representations. The results on SSv2-small [71] and HMDB51 [64] are summarized in Table 9. We can clearly observe that, the performance of our algorithm improves significantly as $\mu$ increases, peaking at $\mu=0.99$ ($60.2\%$ and $85.6\%$). This suggests that a slow and smooth update strategy is essential for maintaining stable and consistent memory representations, effectively mitigating feature fluctuations from individual batches. However, when $\mu$ approaches $1.0$ (e.g., $\mu=0.999$), the performance drops slightly. We attribute this to the memory becoming overly inert, failing to adapt timely to the evolving feature space during training. Notably, at the extreme of $\mu=0$ (i.e., no momentum), the performance degrades sharply ($57.2\%$). This confirms that without the stabilization provided by momentum, the memory bank suffers from rapid changes and noise, hindering effective knowledge transfer.
@@ 4.4.6 Impact of Momentum Coefficient $\mu$ in IKT
Figure 4: The impact of temporal bank size $G$ in IKT on SSv2-small [71] and HMDB51 [64] under the 55-way 11-shot setting (§4.4.7).
Table 10: Impact of module cascading order under the 55-way 11-shot setting on SSv2-small [71] and HMDB51 [64] (see §4.4.8).
     Arrangement    HMDB51    SSv2-small
   IKT $\rightarrow$ ISC    $83.8$    $59.1$
   ISC $\rightarrow$ IKT (Ours)    $\mathbf{85.6}$    $\mathbf{60.2}$
4.4.7 Impact of Bank Size $G$ in IKT
Then we study the influence of our knowledge bank on SSv2-small [71] and HMDB51 [64] in Fig. 4, which determines the capacity for storing temporal representations. As seen from Fig. 4, our algorithm HR2G-shot gains stable improvements (i.e., $59.0\%\rightarrow$$60.2\%$) on SSv2-small [71] as the bank size grows (i.e., $G=50$). This demonstrates that i) there indeed exist some task-shared temporal patterns in the bank, and ii) making use of these task-shared temporal patterns can ensure a sufficient temporal understanding for each task. However, extending the size beyond $50$ (e.g., $G=80$ or $100$) leads to performance saturation or even a slight decline. We attribute this to the inclusion of redundant or less relevant patterns, which may act as noise during the retrieval process. Thus, we set bank size $G$ to $50$ to achieve a better trade-off between accuracy and computation costs.
4.4.8 Module Arrangement Strategy
To validate our “coarse-to-fine” design philosophy, we evaluate the impact of the cascading order between Inter-video Semantic Calibration (ISC) and Inter-task Knowledge Transfer (IKT). As shown in Table 10, the default order (ISC $\rightarrow$ IKT) consistently yields superior performance, surpassing the reverse order by 1.8% and 1.1% on HMDB51 and SSv2-small, respectively. This performance degradation in the reversed order suggests that performing IKT on uncalibrated features may introduce noise or misalignment. These findings confirm that inter-video relation modeling is essential before conducting cross-task knowledge generalization.
@@ 4.4.8 Module Arrangement Strategy
Figure 5: Similarity visualization between query samples ($q_{n}$) and support prototypes ($s_{n}$) with different methods in a meta-test episode from HMDB51 [64] (see §4.5.1). A higher score indicates a greater degree of similarity. The green box indicates correct prediction and the red box indicates incorrect prediction.
@@
Table 11: Complexity analysis for 55-way 11-shot HMDB51 [64] and SSv2-small [71] evaluation. Note that there are 11 query samples per class. We report trainable parameters, GPU memory, and speed for each model. “Acc1” and “Acc2” are the accuarcy on HMDB51 and SSv2-small, respectively (§4.4.9).
@@
CLIP-FSAR 89.589.5M 14.214.2G 36.736.7ms $77.1$ $54.6$
HR2G-shot (ours) $\mathbf{19.9}$M $\mathbf{14.3}$G $\mathbf{39.9}$ms $\mathbf{85.6}$ $\mathbf{60.2}$
4.4.9 Efficiency Analysis
To analyze the effectiveness of our HR2G-shot, we list comparison results with CLIP-FSAR [61] in terms of trainable parameters, GPU memory, and inference speed in Table 11. We choose ViT-B as our visual encoder. Note that different from the full model parameters in Table 4.2, Table 11 reports trainable parameter comparison. As shown in Table 11, our HR2G-shot introduces a lightweight design with only 19.9M trainable parameters, which is substantially lower than the 89.5M used by CLIP-FSAR. Although our approach incurs a marginal overhead in inference time (3.23.2ms delay) and memory consumption (0.10.1G increase) due to the hierarchical reasoning modules, it yields substantial performance gains. Specifically, we achieve improvements of 8.5% and 5.6% in classification accuracy on HMDB51 [64] and SSv2-small [71], respectively. These results validate the effectiveness of our design, confirming that the additional computational cost translates directly into better representation capability.
4.5 Quality Analysis
4.5.1 Similarity Visualization
To qualitatively demonstrate the effectiveness of hierarchical relation modeling in our HR2G-shot, we visualize the predicted similarities between query and support prototypes with different approaches (i.e., CLIP-FSAR [61] and ours) for one task of HMDB51 in Fig. 5. As seen, our HR2G-shot can make more accurate decisions for similar classes in each task compared with CLIP-FSAR. Specifically, for the first query sample in Fig. 5, the incorrect decision obtained by CLIP-FSAR can be rectified by jointly modeling three types of relations. These results further demonstrate that our method can effectively captures task-specific temporal cues by transferring knowledge at both the video level and the task level.
Figure 6: T-SNE feature visualization of six classes learned by baseline and HR2G-shot on Kinetics [70] and HMDB51 [64]4.5.2).
4.5.2 The visualization of Feature Distribution
Fig. 6 visualizes learned features of baseline and our algorithm HR2G-shot with ISC and IKT via the t-SNE tool. We clearly observe that after conducting inter-video and inter-task interactions, learned video features become more compact (intra-class) and better separated (inter-class). As shown in Fig. 6(b), the three classes “Smoke”, “Sit”, and “Pick” become clearly distinguishable from each other after conducting hierarchical relation modeling. The above phenomena suggest that we can gain task-specific temporal features by transferring knowledge at both video-level and task-level, ensuring better feature discrimination.
4.5.3 Attention Visualization of Our HR2G-shot
To qualitatively evaluate the effectiveness of the proposed HR2G-shot, we visualize the attention maps on Kinetics [70] under the 5-way 1-shot setting, as illustrated in Fig. 7. We compare our approach with the baseline CLIP-FSAR [61], which lacks explicit inter-video and inter-task relation modeling. The baseline attention maps are often fragmented—distracted by unrelated background objects, while our HR2G-shot generates more concentrated and precise activations. By integrating the Inter-video Semantic Correlation (ISC) and Inter-task Knowledge Transfer (IKT) modules, our model effectively suppresses task-irrelevant noise and prioritizes action-salient regions. For example, in the “Ice Dancing” sequence, our model consistently anchors its focus on the rhythmic movements of the dancers rather than the static rink environment. These attention Visualization maps provide empirical evidence that hierarchical relation modeling enables the extraction of more discriminative, task-specific temporal features.
Figure 7: Attention visualization of our HR2G-shot on Kinetics in the 5-way 1-shot setting (§4.5.3). Corresponding to the original RGB images, the attention maps of CLIP-FSAR [61] without inter-video and inter-task relation modeling are compared to the attention maps produced by our HR2G-shot.
Figure 8: Generalization performance of our method with different ways and shots on SSv2-full [71] and Kinetics [70]4.6.2).
4.6 Generalization Study
4.6.1 Influence of Different Visual Backbones
Here, we perform a comprehensive evaluation to investigate the generalization capability of our framework across different architectural designs. Specifically, we adopt the ResNet-based CLIP backbone as the visual encoder, while keeping all other algorithmic configurations identical to our default setting. The comparative results against state-of-the-art methods, including CLIP-FSAR [61] and MVP-shot [68], are reported in Table 13. From Table 13, it is evident that our proposed method consistently outperforms existing competitors on the CLIP-RN50 architecture. For instance, under the 1-shot setting, our model surpasses the strong baseline MVP-shot by 1.9%, 1.5%, and 2.0% on SSv2-small [71], UCF101 [65], and HMDB51 [64], respectively. This suggests that our approach is backbone-agnostic and can effectively extract discriminative relational features regardless of the underlying visual encoder.
4.6.2 Results with Various Ways and Shots
We compare HR2G-shot with state-of-the-art methods, including CLIP-FSAR [61], HyRSM [13], OTAM [10], and MTFAN [20], across both $N$-way 1-shot and 5-way $K$-shot scenarios. As shown in Fig. 8, HR2G-shot consistently achieves superior performance across various $N$-way and $K$-shot settings on both SSv2-full [71] and Kinetics [70] datasets. Specifically, on the motion-heavy SSv2-full dataset, our HR2G-shot outperforms the runner-up CLIP-FSAR by 3.2% in the 10-way 1-shot setting (50.1% vs. 46.9%) and maintains a substantial gain of 14.1% over HyRSM in the 5-way 5-shot configuration (74.5% vs. 60.4%). On Kinetics, HR2G-shot achieves peak accuracy of 96.4% in the 5-way 5-shot scenario. These results demonstrate that our hierarchical relation-augmented representations effectively enhance generalization across diverse and demanding few-shot action recognition tasks.